OpenStudy (anonymous):

Write an equation of the line containing the given point and parellel to the given line (6,9); x+8y=3

OpenStudy (amistre64):

parallel lines have the same slope; they lean in the same way...

OpenStudy (amistre64):

and I think they might even have the same equation; but just a different answer; a different "right hand side"... but I would have to take that with a grain of salt.

OpenStudy (anonymous):

Find the slope for that line either by converting into slope intercept or point slope. Then use that same slope and the point you are given and plug it into the point slope formula \[y-y_0 = m(x-x_0)\]

OpenStudy (anonymous):

Where m is the slope of the original line and x0 and y0 are the x and y coordinates of your point

OpenStudy (amistre64):

y = -x/8 + 3/8 y = -x/8 + b (x=6,y=9) 9 = -(6)/8 + b 9 = -3/4 + b 9+3/4 = b 39/4 = b y = -x/8 + 78/8 x+8y = 78 double check: (x=6,y=9) 6+8(9) = 78 6 + 72 = 78 78 = 78 good

OpenStudy (amistre64):

so far, same equation; plug in your x,y and see what the RHS is for a parallel line.... needs more research tho :)

OpenStudy (amistre64):

Does that make sense luv?

OpenStudy (anonymous):

That seems very complicated. \[x+8y=3 \rightarrow y=1/8x + 3/8 \rightarrow slope\ is\ 1/8\] \[New line is y-9 = 1/8(x-6) -> y = x/8 -6/8 +9 \rightarrow 8y -x = 66\]

OpenStudy (anonymous):

Bleh formatting